Tiger Golf Accessories sells golf shoes, gloves, and a laser-guided range-finder that measures distance. Shown below are unitCalculate weighted average unit contribution margin. (Round answer to 2 decimal places e.g. 10.25.) Weighted average unit conCompute the break-even point in units for the company, Break-even point 19100 unitsDetermine the number of units to be sold at the break-even point for each product line. Shoes 5921 pairs of shoes Gloves 7831Determine the number of units to be sold at the break-even point for each product line. Shoes 5921 pairs of shoes Gloves 7831Verify that the mix of sales units to be sold at the break-even point for each product line will generate a zero net income.
Answer #1

Answer

  • Requirement: CVP Income Statement

Sales - Shoes

$621,705

Sales - Gloves

$250,592

Sales - Range Finders

$1,310,260

Total Sales

$2,182,557

Variable Costs

$1,523,989

Contribution margin

$658,568

Fixed Cost

$658,568

Net Income (Loss)

$0

--Working

Sales - Shoes

=5921*105

Sales - Gloves

=7831*32

Sales - Range Finders

=5348*245

Total Sales

$2,182,557

Variable Costs

=2182557-658568

Contribution margin

=19100*34.48

Fixed Cost

658568

Net Income (Loss)

$0